### v1

@@ 6 Second-order optimality on bounded-rank matrices
which is of independent interest schneider2015Lojaconvergence; levin2023remedy. It is shown in section 6.2 that checking second-order optimality over $\mathcal{M}_{\leq r}$ is NP-hard in general. Nevertheless, in section 6.3, we clarify that second-order stationarity remains attainable in certain special cases, through the lens of smooth parameterizations.
@@ 6.2 NP-hardness of verifying second-order optimality
$\mathcal{A}:\mathbb{R}^{m\times n}\to\mathbb{R}^{m\times n}:\,\eta\mapsto(1-\frac{1}{K})\eta-\frac{1}{4}\sum_{(i,j)\in E}(e_{i}e_{j}^{\top}+e_{j}e_{i}^{\top})\eta(e_{i}e_{j}^{\top}+e_{j}e_{i}^{\top}).$
@@ 6.3 Two parameterizations for matrix varieties
Although section 6.2 establishes the NP-hardness of identifying second-order stationary points in general, it remains natural to ask whether finding such points is still possible in certain special cases. To this end, we resort to the technique of smooth parameterization levin2025effect, since the conditions to guarantee $2\!\Rightarrow\!2$” have been developed in section 5.2.
@@
$\mathrm{T}^{2}_{\mathcal{M}_{\leq r}}(X;\eta)=\eta X^{\dagger}\eta+\mathrm{T}_{\mathcal{M}_{r}}(X)=U_{\bot}\dot{C}\varSigma^{-1}\dot{B}V_{\bot}^{\top}+\mathrm{T}_{\mathcal{M}_{r}}(X),$ (6.11)
@@
$\mathrm{\bf Q}_{(L,R),(\dot{L},\dot{R})}(L^{\prime},R^{\prime})\equiv U_{\bot}\dot{C}C^{-\top}B^{-1}\dot{B}V_{\bot}^{\top}\ (\bmod\ \mathrm{T}_{\mathcal{M}_{r}}(X)),$
@@ 7.1 Bouligand tangent cone to the graph
which can be verified by considering the expressions from (2.11) to (2.13).
@@
Without loss of generality, suppose $U_{R}(0)\ U_{R\bot}(0)]=[U_{R}\ U_{R\bot}]$, $\varSigma_{R}(0)=\varSigma_{R}$, and $V_{R}(0)\ V_{R\bot}(0)]=[V_{R}\ V_{R\bot}]$. Since $\operatorname{rank}(R_{1}(t))\equiv k$-$\ell$ and $\operatorname{rank}(\varSigma_{R}(0))=k$-$\ell$, we can find an interval $\left({-\varepsilon,+\varepsilon}\right)$ such that $\varSigma_{R\bot}(t)\equiv 0$, which means $R_{1}(t)=U_{R}(t)\varSigma_{R}(t)V^{\top}_{R}(t)$.
@@
$\mathrm{T}_{\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}}(X,Y)=\left\{{(\eta,\xi)\ \left|\,\begin{array}[]{l}\eta=UAV^{\top}+U_{\bot}B\varSigma V^{\top}+U\varSigma C^{\top}V_{\bot}^{\top}+U_{\bot}KV^{\top}_{\bot}\\ \xi=U_{\bot}DV_{\bot}^{\top}-UB^{\top}RV_{\bot}^{\top}-U_{\bot}RCV^{\top},\\ A\in\mathbb{R}^{s\times s},\,B\in\mathbb{R}^{(m-s)\times s},\,C\in\mathbb{R}^{(n-s)\times s},\\ D\in\mathrm{T}_{\mathbb{R}_{\leq k-r}^{(m-s)\times(n-s)}}(R),\,K\in\mathbb{R}^{(m-s)\times(n-s)}_{\leq r-s},\\ K^{\top}R=0,\,RK^{\top}=0,\\ K^{\top}DV_{R\bot}=0,\ U_{R\bot}^{\top}DK^{\top}=0\end{array}\right.}\right\}\!$.\!\!\!\!\! (7.4)
@@
Next, we turn to show the $\subseteq$” in (7.4). Given any $(\eta,\xi)\in\mathrm{T}_{\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}}(X,Y)$, by definition of the Bouligand tangent cone, it admits sequences $\{t_{k}\}\subseteq\mathbb{R}_{+}$ and $\{(\eta_{k},\xi_{k})\}\subseteq\mathbb{R}^{m\times n}\times\mathbb{R}^{m\times n}$ such that $t_{k}\downarrow 0$, $(\eta_{k},\xi_{k})\to(\eta,\xi)$, and $(X,Y)+t_{k}(\eta_{k},\xi_{k})\in\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}$, or equivalently,
$X+t_{k}\eta_{k}\in\mathcal{M}_{\leq r}\ \ \text{and}\ \ Y+t_{k}\xi_{k}\in\mathrm{N}_{\mathcal{M}_{\leq r}}(X+t_{k}\eta_{k}).$ (7.7)
@@
where $\operatorname{rank}(K)\leq r-s$. Then we are going to determine the relationships between the involved parameters (e.g., $A_{\eta},A_{\xi}$).
@@
$(X+t_{k}\eta_{k})^{\top}(Y+t_{k}\xi_{k})=t_{k}\eta_{k}^{\top}Y+t_{k}X^{\top}\xi_{k}+t_{k}^{2}\eta_{k}^{\top}\xi_{k}=0.$ (7.9)
Dividing the equation by $t_{k}$ and letting $k\to+\infty$ yield $\eta^{\top}Y+X^{\top}\xi=0$, i.e., $VB_{\eta}^{\top}RV_{\bot}+V_{\bot}^{\top}K^{\top}RV_{\bot}^{\top}+V\varSigma A_{\xi}V^{\top}+V\varSigma C_{\xi}V_{\bot}^{\top}=0$, which implies $C_{\xi}=-\varSigma^{-1}B_{\eta}^{\top}R$, $A_{\xi}=0$, and $K^{\top}R=0$. In a similar way, we can derive $B_{\xi}=-RC^{\top}_{\eta}\varSigma^{-1}$ and $RK^{\top}=0$. Substituting these equalities into (7.8) and comparing the expression with (7.4) reduce the task to verifying that $D\in\mathcal{N}(R,K)$ as defined in Lemma 1. To see this, noticing that $U_{\bot}^{\top}\xi_{k}V_{\bot}\to U_{\bot}^{\top}\xi V_{\bot}=D$, and
$U_{\bot}^{\top}(Y+t_{k}\xi_{k})V_{\bot}=R+t_{k}U_{\bot}^{\top}\xi_{k}V_{\bot}\in\mathbb{R}^{(m-s)\times(n-s)}_{\leq k-r}$,
we have $D\in\mathrm{T}_{\mathbb{R}^{(m-s)\times(n-s)}_{\leq k-r}}(R)$. Moreover, pre- and postmultiplying (7.9) by $V_{\bot}V_{\bot}^{\top}$ and $V_{\bot}V_{R\bot}$, respectively, we obtain $V_{\bot}V_{\bot}^{\top}\eta_{k}^{\top}\xi_{k}V_{\bot}V_{R\bot}=0$. Letting $k\to+\infty$ and incorporating the expression of $(\eta,\xi)$ (7.8) reveal that $V_{\bot}(K^{\top}U_{\bot}^{\top}+C^{\top}_{\eta}U^{\top})(UC_{\xi}+U_{\bot}D)V_{R\bot}=0$, which is simplified to $V_{\bot}(K^{\top}D+C_{\eta}^{\top}C_{\xi})V_{R\bot}=0$. Consequently, we use $C_{\xi}V_{R\bot}=-\varSigma^{-1}B_{\eta}^{\top}RV_{R\bot}=0$ to derive $K^{\top}DV_{R\bot}=0$, and similarly, we can also find that $U_{R\bot}^{\top}DK^{\top}=0$, indicating that $D\in\mathcal{N}(R,K)$. Therefore, we have identified all the relationships to conclude that $(\eta,\xi)$ belongs to the set on the right side of (7.4).
@@ 7.3 Mordukhovich normal cone to the graph
  • For all $i_{j}$, $\operatorname{im}[U_{i_{j}}\ \tilde{U}_{i_{j}\bot}]=\operatorname{im}X_{i_{j}}$, $\operatorname{im}[V_{i_{j}}\ \tilde{V}_{i_{j}\bot}]=\operatorname{im}X^{\top}_{i_{j}}$, $\operatorname{im}[U_{i_{j}Y\bot}\ \tilde{U}_{i_{j}Y}]=\operatorname{im}Y_{i_{j}}$, and $\operatorname{im}[V_{i_{j}Y\bot}\ \tilde{V}_{i_{j}Y}]=\operatorname{im}Y^{\top}_{i_{j}}$.
@@
$\upsilon=\left[U^{+}\,\,\breve{U}_{\bot}\,\,U_{Y}^{+}\right]\left[\begin{array}[]{ccc}0&0&B^{\upsilon}\\ 0&Z_{1}&Z_{2}\\ C^{v}&Z_{3}&Z_{4}\end{array}\right]\left[V^{+}\,\,\breve{V}_{\bot}\,\,V_{Y}^{+}\right]^{\top},$ (7.17)
@@
$\hat{\upsilon}_{i}=\left[U^{+}\,\,\breve{U}_{\bot}\,\,U_{Y}^{+}\right]\left[\begin{array}[]{ccc}0&0&B_{i}^{\upsilon}\\ 0&Z_{1}&Z_{2}\\ C^{v}_{i}&Z_{3}&Z_{4}\end{array}\right]\left[V^{+}\,\,\breve{V}_{\bot}\,\,V_{Y}^{+}\right]^{\top},$ (7.20)
@@
We then turn to the “only if” part of the theorem. Suppose that $(\upsilon,\omega)\in\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}(X,Y)$ is generated by the sequence $\{(X_{i},Y_{i})\}_{i\in\mathbb{N}}$ in the sense of  (7.16), and then we will identify that the structure of $(\upsilon,\omega)$ coincides with (7.17). To this end, note that there exists $({\underline{r}},\bar{r})$ such that $s\leq{\underline{r}}\leq r\leq\bar{r}\leq\ell\leq k$ and a subsequence $\{(X_{i_{j}},Y_{i_{j}})\}_{j\in\mathbb{N}}$ satisfying $\operatorname{rank}(X_{i_{j}})={\underline{r}}$ and $\operatorname{rank}(Y_{i_{j}})=k-\bar{r}$ for every $j\in\mathbb{N}$, and the associated subspace sequences (7.15) as constructed in Lemma 3.
@@
$\hat{\upsilon}_{i}$ $=\left[U_{i}^{+}\,\,\breve{U}_{i}\,\,U_{iY}^{+}\right]\left[\begin{matrix}0&0&B_{i}^{\upsilon}\\ 0&Z_{1i}&Z_{2i}\\ C^{v}_{i}&Z_{3i}&Z_{4i}\end{matrix}\right]\left[V_{i}^{+}\,\,\breve{V}_{i}\,\,V_{iY}^{+}\right]^{\top},$
@@
$B^{\upsilon}_{i}=(U^{+}_{i})^{\top}\hat{\upsilon}_{i}V^{+}_{iY},\ C^{\upsilon}_{i}=(U^{+}_{iY})^{\top}\hat{v}_{i}V_{i}^{+},\ Z_{1i}=\breve{U}_{i}^{\top}\hat{\upsilon}_{i}\breve{V}_{i},$
@@
$\upsilon=\left[U^{+}\,\,\breve{U}_{\bot}\,\,U_{Y}^{+}\right]\left[\begin{matrix}0&0&B^{\upsilon}\\ 0&Z_{1}&Z_{2}\\ C^{v}&Z_{3}&Z_{4}\end{matrix}\right]\left[V^{+}\,\,\breve{V}_{\bot}\,\,V_{Y}^{+}\right]^{\top},$ (7.23)
@@ 8.2 Optimality conditions via a relaxation
$\mu\nabla_{X}\mathcal{L}(\tilde{x},\tilde{X})+\mathcal{J}_{X}(\nabla_{X}F)(\tilde{x},\tilde{X})[\delta]+\omega=0,$ (8.2)
$\delta+\upsilon=0,$ (8.3)

### v2

@@ 6 Second-order optimality on bounded-rank matrices
which is of independent interest schneider2015Lojaconvergence; levin2023remedy. It is shown in section 6.2 that checking second-order optimality over $\mathcal{M}_{\leq r}$ is NP-hard in general. Nevertheless, in section 6.3, we clarify that second-order stationarity remains attainable in certain cases, through the lens of smooth parameterizations.
@@ 6.2 NP-hardness of verifying second-order optimality
$\mathcal{A}:\mathbb{R}^{m\times n}\to\mathbb{R}^{m\times n}:\,\eta\mapsto(1-\frac{1}{K-1})\eta-\frac{1}{4}\sum_{(i,j)\in E}(e_{i}e_{j}^{\top}+e_{j}e_{i}^{\top})\eta(e_{i}e_{j}^{\top}+e_{j}e_{i}^{\top}).$
@@ 6.3 Two parameterizations for matrix varieties
Although section 6.2 establishes the NP-hardness of identifying second-order stationary points in general, it remains natural to ask whether finding such points is still possible in special cases. To this end, we resort to the technique of smooth parameterization levin2025effect, since the conditions to guarantee $2\!\Rightarrow\!2$” have been developed in section 5.2.
@@
$\mathrm{T}^{2}_{\mathcal{M}_{\leq r}}(X;\eta)=2\eta X^{\dagger}\eta+\mathrm{T}_{\mathcal{M}_{r}}(X)=2U_{\bot}\dot{C}\varSigma^{-1}\dot{B}V_{\bot}^{\top}+\mathrm{T}_{\mathcal{M}_{r}}(X),$ (6.11)
@@
$\mathrm{\bf Q}_{(L,R),(\dot{L},\dot{R})}(L^{\prime},R^{\prime})\equiv 2U_{\bot}\dot{C}C^{-\top}B^{-1}\dot{B}V_{\bot}^{\top}\ (\bmod\ \mathrm{T}_{\mathcal{M}_{r}}(X)),$
@@ 7.1 Bouligand tangent cone to the graph
which can be verified by considering the conditions $K^{\top}R=0$, $RK^{\top}=0$, and the expressions from (2.11) to (2.13).
@@
Without loss of generality, suppose $U_{R}(0)\ U_{R\bot}(0)]=[U_{R}\ U_{R\bot}]$, $\varSigma_{R}(0)=\varSigma_{R}$, and $V_{R}(0)\ V_{R\bot}(0)]=[V_{R}\ V_{R\bot}]$. Since $\operatorname{rank}(R_{1}(t))\equiv k$-$\ell$ and $\operatorname{rank}(\varSigma_{R}(0))=k$-$\ell$, we can find an $\varepsilon>0$ and the interval $\left({-\varepsilon,+\varepsilon}\right)$ such that $\varSigma_{R\bot}(t)\equiv 0$ for all $t\in(-\varepsilon,+\varepsilon)$, which means $R_{1}(t)=U_{R}(t)\varSigma_{R}(t)V^{\top}_{R}(t)$.
@@
$\mathrm{T}_{\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}}(X,Y)=\left\{{(\eta,\xi)\ \left|\,\begin{array}[]{l}\eta=UAV^{\top}+U_{\bot}B\varSigma V^{\top}+U\varSigma C^{\top}V_{\bot}^{\top}+U_{\bot}KV^{\top}_{\bot}\\ \xi=U_{\bot}DV_{\bot}^{\top}-UB^{\top}RV_{\bot}^{\top}-U_{\bot}RCV^{\top},\\ A\in\mathbb{R}^{s\times s},\,B\in\mathbb{R}^{(m-s)\times s},\,C\in\mathbb{R}^{(n-s)\times s},\\ K\in\mathbb{R}^{(m-s)\times(n-s)}_{\leq r-s},\,K^{\top}R=0,\,RK^{\top}=0,\\ D\in\mathrm{T}_{\mathbb{R}_{\leq k-r}^{(m-s)\times(n-s)}}(R),\\ K^{\top}DV_{R\bot}=0,\ U_{R\bot}^{\top}DK^{\top}=0\end{array}\right.}\right\}\!$.\!\!\!\!\! (7.4)
@@
Next, we turn to show the $\subseteq$” in (7.4). Given any $(\eta,\xi)\in\mathrm{T}_{\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}}(X,Y)$, by definition of the Bouligand tangent cone, it admits sequences $\{t_{i}\}\subseteq\mathbb{R}_{+}$ and $\{(\eta_{i},\xi_{i})\}\subseteq\mathbb{R}^{m\times n}\times\mathbb{R}^{m\times n}$ such that $t_{i}\to 0$, $(\eta_{i},\xi_{i})\to(\eta,\xi)$, and $(X,Y)+t_{i}(\eta_{i},\xi_{i})\in\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}$, or equivalently,
$X+t_{i}\eta_{i}\in\mathcal{M}_{\leq r}\ \ \text{and}\ \ Y+t_{i}\xi_{i}\in\mathrm{N}_{\mathcal{M}_{\leq r}}(X+t_{i}\eta_{i}).$ (7.7)
@@
where $\operatorname{rank}(K)\leq r-s$. Then we will determine the relationships between the involved parameters (e.g., $A_{\eta},A_{\xi}$).
@@
$(X+t_{i}\eta_{i})^{\top}(Y+t_{i}\xi_{i})=t_{i}\eta_{i}^{\top}Y+t_{i}X^{\top}\xi_{i}+t_{i}^{2}\eta_{i}^{\top}\xi_{i}=0.$ (7.9)
Dividing the equation by $t_{i}$ and letting $i\to+\infty$ yield $\eta^{\top}Y+X^{\top}\xi=0$, i.e., $VB_{\eta}^{\top}RV_{\bot}^{\top}+V_{\bot}K^{\top}RV_{\bot}^{\top}+V\varSigma A_{\xi}V^{\top}+V\varSigma C_{\xi}V_{\bot}^{\top}=0$, which implies $C_{\xi}=-\varSigma^{-1}B_{\eta}^{\top}R$, $A_{\xi}=0$, and $K^{\top}R=0$. In a similar way, we can derive $B_{\xi}=-RC^{\top}_{\eta}\varSigma^{-1}$ and $RK^{\top}=0$. Substituting these equalities into (7.8) and comparing the expression with (7.4) reduce the task to verifying that $D\in\mathcal{N}(R,K)$ as defined in Lemma 1. To see this, noticing that $U_{\bot}^{\top}\xi_{i}V_{\bot}\to U_{\bot}^{\top}\xi V_{\bot}=D$, and
$U_{\bot}^{\top}(Y+t_{i}\xi_{i})V_{\bot}=R+t_{i}U_{\bot}^{\top}\xi_{i}V_{\bot}\in\mathbb{R}^{(m-s)\times(n-s)}_{\leq k-r}$,
we have $D\in\mathrm{T}_{\mathbb{R}^{(m-s)\times(n-s)}_{\leq k-r}}(R)$. Moreover, pre- and postmultiplying (7.9) by $V_{\bot}V_{\bot}^{\top}$ and $V_{\bot}V_{R\bot}$, respectively, we obtain $V_{\bot}V_{\bot}^{\top}\eta_{i}^{\top}\xi_{i}V_{\bot}V_{R\bot}=0$. Letting $i\to+\infty$ and incorporating the expression of $(\eta,\xi)$ (7.8) reveal that $V_{\bot}(K^{\top}U_{\bot}^{\top}+C^{\top}_{\eta}U^{\top})(UC_{\xi}+U_{\bot}D)V_{R\bot}=0$, which is simplified to $V_{\bot}(K^{\top}D+C_{\eta}^{\top}C_{\xi})V_{R\bot}=0$. Consequently, we use $C_{\xi}V_{R\bot}=-\varSigma^{-1}B_{\eta}^{\top}RV_{R\bot}=0$ to derive $K^{\top}DV_{R\bot}=0$, and similarly, we can also find that $U_{R\bot}^{\top}DK^{\top}=0$, indicating that $D\in\mathcal{N}(R,K)$. Therefore, we have identified all the relationships to conclude that $(\eta,\xi)$ belongs to the set on the right side of (7.4).
@@ 7.3 Mordukhovich normal cone to the graph
  • For all $i_{j}$, $\operatorname{im}[U_{i_{j}}\ \tilde{U}_{i_{j}\bot}]=\operatorname{im}X_{i_{j}}$, $\operatorname{im}[V_{i_{j}}\ \tilde{V}_{i_{j}\bot}]=\operatorname{im}X^{\top}_{i_{j}}$, $\operatorname{im}[\tilde{U}_{i_{j}Y\bot}\ U_{i_{j}Y}]=\operatorname{im}Y_{i_{j}}$, and $\operatorname{im}[\tilde{V}_{i_{j}Y\bot}\ {V}_{i_{j}Y}]=\operatorname{im}Y^{\top}_{i_{j}}$.
@@
$\upsilon=\left[U^{+}\,\,\breve{U}_{\bot}\,\,U_{Y}^{+}\right]\left[\begin{array}[]{ccc}0&0&B^{\upsilon}\\ 0&Z_{1}&Z_{2}\\ C^{\upsilon}&Z_{3}&Z_{4}\end{array}\right]\left[V^{+}\,\,\breve{V}_{\bot}\,\,V_{Y}^{+}\right]^{\top},$ (7.17)
@@
$\hat{\upsilon}_{i}=\left[U^{+}\,\,\breve{U}_{\bot}\,\,U_{Y}^{+}\right]\left[\begin{array}[]{ccc}0&0&B_{i}^{\upsilon}\\ 0&Z_{1}&Z_{2}\\ C^{\upsilon}_{i}&Z_{3}&Z_{4}\end{array}\right]\left[V^{+}\,\,\breve{V}_{\bot}\,\,V_{Y}^{+}\right]^{\top},$ (7.20)
@@
We then turn to the “only if” part of the theorem. Suppose that $(\upsilon,\omega)\in\mathrm{N}_{\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}}(X,Y)$ is generated by the sequence $\{(X_{i},Y_{i})\}_{i\in\mathbb{N}}$ in the sense of  (7.16), and then we will identify that the structure of $(\upsilon,\omega)$ coincides with (7.17). To this end, note that there exists $({\underline{r}},\bar{r})$ such that $s\leq{\underline{r}}\leq r\leq\bar{r}\leq\ell\leq k$ and a subsequence $\{(X_{i_{j}},Y_{i_{j}})\}_{j\in\mathbb{N}}$ satisfying $\operatorname{rank}(X_{i_{j}})={\underline{r}}$ and $\operatorname{rank}(Y_{i_{j}})=k-\bar{r}$ for every $j\in\mathbb{N}$, and the associated subspace sequences (7.15) as constructed in Lemma 3.
@@
$\hat{\upsilon}_{i}$ $=\left[U_{i}^{+}\,\,\breve{U}_{i}\,\,U_{iY}^{+}\right]\left[\begin{matrix}0&0&B_{i}^{\upsilon}\\ 0&Z_{1i}&Z_{2i}\\ C^{\upsilon}_{i}&Z_{3i}&Z_{4i}\end{matrix}\right]\left[V_{i}^{+}\,\,\breve{V}_{i}\,\,V_{iY}^{+}\right]^{\top},$
@@
$B^{\upsilon}_{i}=(U^{+}_{i})^{\top}\hat{\upsilon}_{i}V^{+}_{iY},\ C^{\upsilon}_{i}=(U^{+}_{iY})^{\top}\hat{\upsilon}_{i}V_{i}^{+},\ Z_{1i}=\breve{U}_{i}^{\top}\hat{\upsilon}_{i}\breve{V}_{i},$
@@
$\upsilon=\left[U^{+}\,\,\breve{U}_{\bot}\,\,U_{Y}^{+}\right]\left[\begin{matrix}0&0&B^{\upsilon}\\ 0&Z_{1}&Z_{2}\\ C^{\upsilon}&Z_{3}&Z_{4}\end{matrix}\right]\left[V^{+}\,\,\breve{V}_{\bot}\,\,V_{Y}^{+}\right]^{\top},$ (7.23)
@@ 8.2 Optimality conditions via a relaxation
$\mu\nabla_{X}\mathcal{L}(\tilde{x},\tilde{X})+\mathcal{J}_{X}(\nabla_{X}F)(\tilde{x},\tilde{X})[\delta]+\upsilon=0,$ (8.2)
$\delta+\omega=0,$ (8.3)